OpenStudy (anonymous):

If y=2/3x+4, find y if a. x=0 b. x=3 c. x=9 d. x= -9 e. x=10 f. x=1/2

OpenStudy (anonymous):

You just need to substitute each value of x in for the x in the equation and see what it equals. So for x=0, you'd simplify this: \[\frac{ 2 }{ 3 }(0)+4\]

OpenStudy (anonymous):

can someone help me

OpenStudy (anonymous):

il show u how to do 2, just do it the way fr the rest a) x=0 we have y=2/3*x + 4 okay put x=0, what do we have Y=2/3*0+4=4 b)x=3 y=2/3*3+4=2+4=6 just do it the same way for the rest:)

OpenStudy (anonymous):

thanks I really appreciate it